# Gen-C: Populating Virtual Worlds with Generative Crowds

### v3

@@ 7.3 Rationale for a Dedicated Learned Generator
Scalability, Validity, and Cost. We compare vanilla LLM generation against our learned generator (Gen-C) on University Campus under increasing agent counts $A_{count}\in[20,160]$. For consistency we proportionally scale the requested event range and the LLM’s max response tokens, and evaluate $50$ seeds per setting while reusing the same $S_{in}$ and a fixed environment designed per seed. The LLM uses the $Q2$ query (Section 4.1), updating agents/events/tokens per setting; Gen-C samples subgraphs from its prior and composes them until the target $A_{count}$ is reached. For each run we report four metrics: (i) action-sequence entropy $H_{\text{seq}}$ (normalized entropy over per-agent action sequences), (ii) inference time (LLM: end-to-end completion; Gen-C: sampling + validation over composed subgraphs), (iii) token response usage (LLM only), and (iv) validity (LLM only) as the drop rate, i.e., the fraction of seeds whose realized agent count falls outside a tolerance $[0.8,1.2]$ of the target. In Fig. 5, we report means for metrics with $95\%$ confidence intervals.
As $A_{count}$ grows, the LLM shows a monotonic decrease in $H_{\text{seq}}$ (agents follow increasingly similar plans), a rising drop rate (more failures to meet the target $A_{count}$, and sharply increasing token usage and latency. In contrast, Gen-C maintains stable action sequence diversity and low, slowly increasing inference time. We also tested smaller LLMs (gpt-4.1-mini and gpt-4.1-nano), which exhibited high drop rates ($0.32$ and $0.94$, respectively) even at baseline $A_{count}=20$. Overall, we show that a dedicated learned generator preserves diversity and reliability at scale, whereas prompt-based LLM generation degrades in both quality and efficiency under the same conditioning, revealing the need of a learned generator.
@@
Prompt Engineering. LLM-based crowd scenario generation currently relies on hand-crafted rules and constraints, creating a heavy configuration burden for non-experts. We employ a schema-guided query enforcing hard constraints (e.g., agent/event limits, diversity, no overlaps). Achieving validity requires (i) detailed instructions with few-shot examples, (ii) a custom JSON schema with post-validators, (iii) tuning of decoding parameters, and (iv) stepwise prompting with retries. These settings are brittle: small changes in scene scale and context often demand re-tuning, making the process error-prone; prompt details are in the supplementary material. In contrast, our learned generator needs only a short description and a few intuitive scalars (#agents, #nodes, #events, action-frequency priors). Without prompt engineering, it sustains high validity, fidelity, and coverage with predictable latency and cost—enabling novice users to easily populate virtual scenes with diverse crowds.
@@ 7.4 Quantitative Results
Figure 6: Graph Structure and Feature reconstruction quality from sampling using the conditional prior on unseen data.
@@
As shown in Fig. 6, we compute the KLD between ground-truth and reconstructed graph distributions; lower values indicate better alignment. VGAE-S accurately recovers structural statistics, while VGAE-F achieves strong feature reconstruction with label distributions close to the ground truth. Performance is consistent across datasets with different dynamics. For example, in the Train Station dataset “wait” is more common and “meet” less so compared to University Campus, reflecting context-specific behaviors such as travelers waiting for trains versus students meeting on campus. Similarly, “service station” location appears more frequently in Train Station, corresponding to information booths or arrival/departure boards, whereas outdoor areas are more prevalent in the University Campus scenario. Additionally, “object interact” and “discuss” stand out in both datasets; the former capturing context-specific actions (e.g., eating in a campus cafeteria or grabbing a ticket at a station), and the latter reflecting a common social behavior.
@@
Latent Space Analysis. We assess how closely graphs sampled from the model prior match the training distribution using a fixed evaluator: encoders trained only on training data. For each theme (University Campus & Train Station), we embed the training set $T$ and predicted set $P$ into latent spaces $Z_{T}$ and $Z_{P}$, derived from encoder posterior means $\mu$ for structure and features. We compute the Fréchet Inception Distance (FID) [66] and Maximum Mean Discrepancy (MMD) [67]: FID compares first- and second-order moments of the latent distributions, while MMD captures higher-order and local differences via a kernel-based similarity. Together, they provide complementary measures of real-generated space alignment.
@@
$\mathrm{FID}$ $=\lVert\mu_{T}-\mu_{P}\rVert_{2}^{2}+\mathrm{Tr}\!\left(\Sigma_{T}+\Sigma_{P}-2(\Sigma_{T}\Sigma_{P})^{1/2}\right),$ (9)
$\mathrm{MMD}^{2}$ $=\mathbb{E}_{TT}[k_{\gamma}(x,x^{\prime})]+\mathbb{E}_{PP}[k_{\gamma}(y,y^{\prime})]-2\,\mathbb{E}_{TP}[k_{\gamma}(x,y)],$
where $k_{\gamma}(x,y)=\exp(-\gamma\|x-y\|_{2}^{2})$.
Table 2: Evaluator-space metrics for real vs. sampled graphs.
@@
Uni. Campus Structure 1.950 0.078 2.921 0.099
Features 9.286 0.032 17.578 0.091
Train Station Structure 0.534 0.058 3.979 0.076
Features 7.674 0.030 16.004 0.109
As shown in Table 2, FID and MMD remain low (FID: 0.53–9.29; MMD: 0.030–0.078) across both themes, with Train Station aligning closer to the training distribution than University Campus. The structure model yields lower FID and feature model lower MMD: structure latents match global moments, while feature latents preserve local neighborhoods. We also run cross-domain checks by mixing model/data: the model trained on University Campus generates samples on Train Station and vice versa. As reported in Table 2 (Mix), both FID and MMD degrade, reflecting reduced alignment across domains; this confirms that Gen-C learns appropriate latent spaces, capturing unique structure and dynamics of each domain.
Figure 7: 2K True/Predicted latent vectors for Train Station.
Moreover, in Fig. 7, we plot 2K true and predicted latent vectors via PCA for VGAE-S (Structure) and VGAE-F (Features) on Train Station. In-domain, true and predicted latents largely co-locate. We also show cross-domain “mix-and-match” (Mix) projections, which exhibit reduced overlap and clear domain shifts, consistent with degraded FID/MMD in Table 2. The mixed-structure space retains some overlap, as similar structures span themes, whereas the feature latent space shows a strong mismatch, indicating that node features are tightly coupled to input context. Overall, both metrics and embeddings suggest our generator produces crowd scenario graphs well aligned with the training distribution in evaluator space.
@@ 7.5 Qualitative Results
In Fig. 8, we show rendered results across both datasets. Gen-C reproduces sequences of behaviors observed in real environments, including agents queuing for food or coffee, reading announcements, waiting for trains, chatting over lunch, and wandering through open areas. These examples highlight the model’s ability to generate heterogeneous yet coherent multi-agent actions and interactions. Additional qualitative results and longer sequences are provided in the supplementary video.
Figure 8: Generated render results from our framework.
@@ 8 Discussion and Future Work
We introduce Gen-C, a framework for populating virtual scenes with multiple agents that perform diverse, high-level actions and interactions from a short textual prompt. Gen-C couples a schema-guided LLM with two learned latent spaces (graph structure & features) to synthesize crowd scenario graphs that drive agents behavior. In contrast to traditional crowd systems focused on local navigation (e.g., goal seeking and collision avoidance), Gen-C targets the semantic layer—modeling activities, locations, and social interactions—to generate plausible, varied, and scalable scenarios.
Our current implementation does not support modeling long-term agent intentions, agents cannot switch actions mid-execution, and action durations are sampled from manually defined distributions. Additionally, behaviors are constrained to a predefined list of high-level actions, requiring retraining for new actions.
For future work, we aim to equip agents with a memory module, ensuring a more natural progression of actions. We would also like to combine Gen-C with a crowd simulator, allowing high-level behaviors to control low-level navigation policies, creating more dynamic and plausible simulations. Furthermore, we plan to explore combining latent spaces from different datasets, such as indoor/outdoor scenarios, using techniques such as distillation to enable lifelong learning and dynamic simulation adaptability. We envision that these will allow us to populate virtual worlds with diverse human-like characters, paving the road towards automatic population of virtual worlds.

### v4

@@ 7.2 Rationale for a Dedicated Learned Generator
Scalability, Validity, and Cost. We compare vanilla LLM generation against our learned generator (Gen-C) on University Campus under increasing agent counts $A_{count}\in[20,160]$. For consistency we proportionally scale the requested event range and the LLM’s max response tokens, and evaluate $50$ seeds per setting while reusing the same $S_{in}$ and a fixed environment designed per seed. The LLM uses the $Q2$ query (Section 4.1), updating agents/events/tokens per setting; Gen-C samples subgraphs from its prior and composes them until the target $A_{count}$ is reached. For each run we report four metrics: (i) action-sequence entropy $H_{\text{seq}}$ (normalized entropy over per-agent action sequences), (ii) inference time (LLM: end-to-end completion; Gen-C: sampling + validation over composed subgraphs), (iii) token response usage (LLM only), and (iv) validity (LLM only) as the drop rate, i.e., the fraction of seeds whose realized agent count falls outside a tolerance $[0.8,1.2]$ of the target. In Figure 6, we report means for metrics with $95\%$ confidence intervals.
As $A_{count}$ grows, the LLM shows a monotonic decrease in $H_{\text{seq}}$ (agents follow increasingly similar plans), a rising drop rate (more failures to meet the target $A_{count}$, and sharply increasing token usage and latency. In contrast, Gen-C maintains stable action sequence diversity and low, slowly increasing inference time. Smaller LLMs (gpt-4.1-mini and gpt-4.1-nano), show high drop rates ($0.32$ and $0.94$) even at baseline $A_{count}=20$. Overall, we show that a dedicated learned generator preserves diversity and reliability at scale, whereas prompt-based LLM generation degrades in both quality and efficiency under the same conditioning, revealing the need of a learned generator.
@@
Figure 7: Graph structure and feature reconstruction quality on unseen data using samples from the conditional prior. For structure, we report various graph metrics, while for features we compare label frequencies.
Prompt Engineering. LLM-based crowd scenario generation currently relies on hand-crafted rules and constraints, creating a heavy configuration burden for non-experts. We employ a schema-guided query enforcing hard constraints (e.g., agent/event limits, diversity, no overlaps). Achieving validity requires (i) detailed instructions with few-shot examples, (ii) a custom JSON schema with post-validators, (iii) tuning of decoding parameters, and (iv) stepwise prompting with retries. These settings are brittle: small changes in scene scale and context often demand re-tuning, making the process error-prone; prompt details are in the supplementary material. In contrast, our learned generator needs only a short description and a few intuitive scalars (#agents, #nodes, #events, action-frequency priors). Without prompt engineering, it sustains high validity, fidelity, and coverage with predictable latency and cost, enabling novice users to easily populate scenes with diverse crowds.
@@ 7.3 Quantitative Results
As shown in Figure 7, we compute the KLD between ground-truth and reconstructed graph distributions; lower values indicate better alignment. VGAE-S accurately recovers structural statistics, while VGAE-F achieves strong feature reconstruction with label distributions close to the ground truth. Performance is consistent across datasets with different dynamics. For example, in the Train Station dataset “wait” is more common and “meet” less so compared to University Campus, reflecting context-specific behaviors such as travelers waiting for trains versus students meeting on campus. Similarly, “service station” location appears more frequently in Train Station, corresponding to information booths or arrival/departure boards, whereas outdoor areas are more prevalent in the University Campus scenario. Additionally, “object interact” and “discuss” stand out in both datasets; the former capturing context-specific actions (e.g., eating in a campus cafeteria or grabbing a ticket at a station), and the latter reflecting a common social behavior.
Table 3: Top-4 action-location pairs predicted by LLM and Gen-C per theme.
University Campus
LLM sit@furniture =0.08=0.08, enter/exit@entrance =0.07=0.07, wander@outdoor area =0.07=0.07, discuss@furniture =0.05=0.05
Gen-C sit@furniture =0.06=0.06, enter/exit@entrance =0.05=0.05, wander@outdoor area =0.05=0.05, wave at@outdoor area =0.04=0.04
Train Station
LLM enter/exit@entrance =0.08=0.08, sit@furniture =0.08=0.08, look at@item =0.07=0.07, object interact@service station =0.06=0.06
Gen-C enter/exit@entrance =0.07=0.07, look at@item =0.06=0.06, sit@furniture =0.06=0.06, queue@service station =0.05=0.05
Additionally, scenario differences arise not only from action frequencies but also from their joint patterns with locations. For each theme, we analyze the joint distribution of actions and locations by aggregating all observed action@location pairs and computing their empirical frequencies for both LLM generations and Gen-C predictions. We select the top-k pairs, normalize their counts by the total occurrences, and report the top-4 per theme in Table 3. The results show that: (i) our model preserves the dominant interaction patterns, and (ii) the prevalent pairs vary meaningfully across themes, reflecting contextual differences, while maintaining behavioral diversity.
@@
Latent Space Analysis. We assess how closely graphs sampled from the model prior match the training distribution using a fixed evaluator: encoders trained only on training data. For each theme (University Campus & Train Station), we embed the training set $T$ and predicted set $P$ into latent spaces $Z_{T}$ and $Z_{P}$, derived from encoder posterior means $\mu$ for structure and features. We compute the Fréchet Inception Distance (FID) [13] and Maximum Mean Discrepancy (MMD) [11]: FID compares first- and second-order moments of the latent distributions, while MMD captures higher-order and local differences via a kernel-based similarity. Taken together, they provide complementary measures of alignment between the real and generated spaces, and are defined as follows:
@@
$\begin{gathered}\mathrm{FID}=\lVert\mu_{T}-\mu_{P}\rVert_{2}^{2}+\mathrm{Tr}\!\left(\Sigma_{T}+\Sigma_{P}-2(\Sigma_{T}\Sigma_{P})^{1/2}\right),\\ \mathrm{MMD}^{2}=\mathbb{E}_{TT}[k_{\gamma}(x,x^{\prime})]+\mathbb{E}_{PP}[k_{\gamma}(y,y^{\prime})]-2\,\mathbb{E}_{TP}[k_{\gamma}(x,y)],\\ k_{\gamma}(x,y)=\exp(-\gamma\|x-y\|_{2}^{2}).\end{gathered}$ (6)
As shown in Table 4, FID and MMD remain low (FID: 0.53-9.29; MMD: 0.030-0.078) across both themes, with Train Station (TS) aligning closer to the training distribution than University Campus (UC). The structure model (S) yields lower FID and feature model (F) lower MMD: structure latents match global moments, while feature latents preserve local neighborhoods. We also run cross-domain checks by mixing model/data. Specifically, we use the model trained on University Campus to generate samples on Train Station and vice versa. We show that both FID (Mix) and MMD (Mix) degrade, reflecting reduced alignment across domains; this confirms that Gen-C learns appropriate latent spaces, capturing unique structure and dynamics of each domain.
Table 4: Latent space metrics.
@@
UC S 1.9503 0.0779 2.9205 0.0986
F 9.2861 0.0319 17.5777 0.0913
TS S 0.5341 0.0575 3.9793 0.0757
F 7.6735 0.0296 16.0035 0.1091
Moreover, in Figure 8, we plot 2K true and predicted latent vectors via PCA for VGAE-S (Structure) and VGAE-F (Features) on Train Station theme. In-domain, true and predicted latents largely co-locate. We also show cross-domain “mix-and-match” (Mix) projections, which exhibit reduced overlap and clear domain shifts, consistent with degraded FID/MMD in Table 4. The mixed-structure space retains some overlap, as similar structures span themes, whereas the feature latent space shows a strong mismatch, clearly indicating that node features are tightly coupled to the input context. Overall, both metrics and embeddings suggest that our learned generator produces crowd scenario graphs that are well aligned with the training distribution in the evaluator space.
Figure 8: 2K latent vectors for Train Station.
@@ 7.4 Qualitative Results
Figure 9: (Left) Comparison of action probabilities for the specific user study scenarios. (Right) Comparison of action probabilities for the general user study scenarios.
User Study. To assess the semantic plausibility of the generated crowd behaviors and their alignment with real-world conditions, we conducted a user study ($N=29$) comparing Gen-C predicted action distributions against human expectations. Participants were recruited through the authors’ professional networks and the Prolific platform [47] and completed the study via an anonymous online questionnaire. Expertise in crowd dynamics varied across respondents, with scores centered near the midpoint of the 7-point Likert scale ($M=3.50$, $SD=1.44$), and participant ages ranged from 18 to 64 years. For each of the two themes, participants reviewed seven textual descriptions (14 in total) of specific scenarios ($S_{in}$) and were asked to select the top-5 most likely actions from $\mathbf{\mathcal{A}ct}$, ranking them from most to least common; actions were presented in a random order to avoid bias. Additionally, to evaluate the model’s ability to capture broader environmental “rules”, we collected responses for four abstract, context-free queries (e.g., “You are standing on a train platform. What actions do you expect people around you to be doing?”). Details of the user study and its questions appear in the supplementary material.
We prioritized a text-based evaluation to rigorously isolate the model’s behavioral logic from visual artifacts. In video-based assessments, particularly in multi-agent settings, user perception is often strongly influenced by rendering quality, animation smoothness, and the fidelity of motion and interactions. By removing these confounding factors, the selected assessment avoids the substantial overhead of constructing scenario-specific 3D environments and enables a focused evaluation of the semantic plausibility of the generated scenarios. Figure 9 shows the alignment of actions between Gen-C predictions and user responses for both the (a) specific and (b) general scenarios. Specifically, for (a), we run our model on the 14 specific scenarios used in the study (20 runs per $S_{in}$) and aggregate the resulting action probabilities. The model demonstrates robust alignment with human expectations across both themes, achieving low Jensen-Shannon Divergence scores ($JSD\approx 0.24$). For (b), we compare the model’s global behavior (aggregated across the full test set) against human expectations for abstract, context-free queries. Notably, the results highlight the model’s ability to capture distinct environmental “rules”: it shows stronger convergence in the structured, rule-bound Train Station ($JSD=0.172$) compared to the unstructured, high-entropy University Campus ($JSD=0.308$), correctly reflecting the wider behavioral variance inherent to open social spaces; additional details are provided in the supplementary material.
Comparison with Baseline Crowd Models. As discussed, existing crowd models primarily address local navigation and collision avoidance, and thus operate at a different layer of the simulation stack than our approach. Recent text-guided crowd synthesis methods [17] focus on group-level, goal-driven navigation, while frameworks such as GREIL-Crowds [5] and CEDRL [42] introduce some degree of behavioral variation but their behaviors are limited to coarse categories like moving, idling, or loosely grouped motion. The interactions between agents typically emerge implicitly from low-level navigation dynamics, while interactions with the environment are minimal or absent. Moreover, these methods are designed to generate short-horizon trajectories rather than explicit, structured behavior sequences, and cannot be conditioned on specific scenario-level descriptions beyond the general environment type (e.g., a university campus or a shopping street). In contrast, Gen-C targets the high-level decision-making and planning layer of a crowd simulator, explicitly generating scenario-conditioned sequences of semantic actions and interactions that define agent-agent and agent-environment behavior over time. Thus, Gen-C is intended to complement existing crowd simulation approaches. As a result, our comparison with baseline models should be interpreted as a qualitative and coarse-grained analysis of behavioral diversity rather than a direct performance benchmark. We refer readers to the supplementary video for simulations produced by both the baseline models and our method.
Figure 10: Our model can populate environments with rich, engaging crowd behaviors. Agents form queues for food or coffee, read announcements, wait for trains, converse over lunch, wander around, and more.
Rendered Results. In Figure 10, we present rendered results across both themes; scenarios are simulated and visualized in Unity [61], with character animations manually assigned based on the high-level action each agent is performing. Gen-C generates semantically plausible behaviors, including agents queuing for food or coffee, reading announcements, waiting for trains, chatting over lunch, meeting friends, wandering through open areas, and more. These examples demonstrate the model’s ability to produce heterogeneous yet coherent multi-agent actions and interactions. Additional qualitative results are included in the supplementary video.
@@ 8 Discussion and Future Work
We introduce Gen-C, a framework for populating virtual scenes with multiple agents that perform diverse, high-level actions and interactions from a short textual prompt. Our framework couples a schema-guided LLM with two learned latent spaces (graph structure & features) to synthesize crowd scenario graphs that drive agents behavior. In contrast to traditional crowd systems focused on local navigation, Gen-C targets the semantic layer, modeling activities, locations, and social interactions, to generate plausible, varied, and scalable scenarios.
Our current implementation does not support modeling long-term agent intentions, agents cannot switch actions mid-execution, and action durations are sampled from manually defined distributions. Additionally, behaviors are constrained to a predefined list of high-level actions, constraining behavioral expressiveness and requiring retraining for new actions; environmental context is also limited to a simple positional map with constraint location types. These limitations reflect our focus on high-level planning rather than fine-grained control, and highlight several opportunities for future research. Implementing memory or belief states could enable agents to reason over past interactions and produce more coherent long-term behavior. Incorporating geometry-aware and physical feasibility constraints like density, traversability maps, and proximity penalties would ensure spatially consistent behaviors, while expanding actions via hierarchical taxonomies and hybrid datasets would boost adaptability. We would also like to integrate Gen-C with existing crowd simulators, enabling high‑level plans to guide low‑level navigation policies and thereby bridging semantic planning with physical motion. Finally, we plan to explore combining latent spaces learned from heterogeneous datasets (e.g., indoor and outdoor scenes) using techniques such as distillation or continual learning, enabling adaptive and extensible crowd generation across diverse settings. We view our work as a step toward scalable, data-driven population of virtual worlds with rich, human-like behavior.